# Leveraging Discourse Structure for Extractive Meeting Summarization

### v2

@@ Abstract
We introduce an extractive summarization system for meetings that leverages discourse structure to better identify salient information from complex multi-party discussions. Using discourse graphs to represent semantic relations between the contents of utterances in a meeting, we train a GNN-based node classification model to select the most important utterances, which are then combined to create an extractive summary. Experimental results on AMI and ICSI demonstrate that our approach surpasses existing text-based and graph-based extractive summarization systems, as measured by both classification and summarization metrics. Additionally, we conduct ablation studies on discourse structure and relation type to provide insights for future NLP applications leveraging discourse analysis theory.
@@ 1 Introduction
In recent years, the task of meeting summarization has garnered significant attention Rennard et al. (2023b) due in part to the rising adoption of videoconferencing Kost (2020) and the accumulated amount of meeting recordings, highlighting the need for automatic, novel and efficient processing methods. Despite recent advancements in natural language processing, the challenge of generating concise and coherent summaries from lengthy, multi-party meeting transcriptions remains. Large language models suffer from the middle curse (Liu et al., 2024) and struggle to use information in the middle of their context window, particularly for summarization (Ravaut et al., 2023).
@@
Our approach involves a novel combination of Graph Neural Networks (GNN) and graph-based representations of discourse in which each node in a graph represents the content of an individual utterance and each edge represents a semantic relation between two utterances whose nature is specified by a label, e.g., Explanation, Correction, Question-Answer Pair or Acknowledgment. The task of extractive summarization is then couched as one of binary node-classification in which nodes are determined to be important or not, and the content of the nodes judged to be important is what determines the final extractive summary. This approach, in contrast to the generation at the whole graph-level (Feng et al., 2020), allows us to gain fine-grained insight into the interaction between the importance of an utterance in a conversation and its role in the overall discourse. A further advantage of a graph-based approach is that GNN, whose attention is focused only on the neighbors of a given node, do not suffer from the context length restrictions imposed by the pairwise attentional computations of transformer-based models. This means that our extractive summarization approach can take an entire long meeting transcript as input without special treatment.
To validate our discourse-structure approach, we conduct extensive experiments on the AMI (Mccowan et al., 2005) and ICSI (Janin et al., 2003) corpora, comparing our method against diverse extractive summarization systems. The results demonstrate that our approach outperforms traditional techniques across evaluation metrics for both classification and summarization, including F1 score, ROUGE and BERTScore. Additionally, we present a detailed analysis of the impact of various graph construction strategies on summarization quality, offering insights into the mechanisms through which discourse structure influences content selection. Our study not only provides a novel methodological contribution to the field of automatic meeting summarization but also sheds light on the underlying discourse processes that shape effective summaries. Finally, discourse structure has seldom been explored in the literature for its potential utility in downstream tasks. Our work is among the few studies that focus on this effort, especially for dialogues.
@@ 2 Related work
Most work focuses on abstractive summarization and tends to be organized in three interconnected streams (Rennard et al., 2023b). The first focuses on enhancing transcripts with additional information—such as annotations for dialogue acts (Goo and Chen, 2018), discourse structure (Feng et al., 2020) or visual cues (Liu et al., 2019)—that is assumed to be relevant for abstracting important content from a transcript. While adding information does show improvement, few papers actually quantify how much information is gained from adding linguistical features; Indeed, Goo and Chen (2018) worked on a version of AMI that lacked the current version of abstractive summaries, making it difficult to draw conclusions that are comparable to more recent works. Additionally, Feng et al. (2020) used a global node to represent the discourse graph, aggregating all features into a single representative vector sent to an LSTM-based pointer decoder network for summarization. This approach is not ideal for evaluating the role of discourse structure in enhancing abstractive summaries. The second stream aims to transform and compress meeting transcripts in order to produce cleaner and more condensed intermediate documents that can then be passed to downstream generation modules (Krishna et al., 2021; Oya et al., 2014). The third stream is dedicated to the development of general generation frameworks, such as DialogLM (Zhong et al., 2022), that are specifically designed to handle dialogue content. While we focus on extractive summarization, in exploiting discourse structure, our approach overlaps with work in the first stream above and lays groundwork for the second stream by providing a means of compressing meeting transcripts.
@@
Two primary theoretical frameworks have been developed to analyze complete discourse structures in the form of graphs, namely Rhetorical Structure Theory (RST; Mann and Thompson, 1987) and Segmented Discourse Representation Theory (SDRT; Asher, 1993; Lascarides and Asher, 2008). The framework of RST has been successfully applied for extractive summarization of newspaper articles (Xu et al., 2020; Liu and Chen, 2019). However most work on discourse graphs has concentrated primarily on producing high-quality graphs (Shi and Huang, 2019; Liu and Chen, 2021; Bennis et al., 2023), with less emphasis on their applications to downstream tasks such as summarization, and only SDRT has been applied to build discourse structures for multi-party conversation (Asher et al., 2016), which is needed for meeting summarization.
A major limitation for research on discourse structure is of course the intense effort that goes in to annotating datasets for discourse structure. There are only two readily available discourse-annotated corpora for multiparty conversation: STAC (Asher et al., 2016), a corpus of chats from an online version of the game Settlers of Catan, and Molweni (Li et al., 2020), an annotated version of the Ubuntu Chat Corpus. Both corpora are annotated in the style of SDRT.
@@ 3 Preliminaries
Due to its appropriateness for multiparty dialogue (Asher et al., 2016) and to the fact that it is the framework adopted for the STAC and Molweni corpora, we focus on SDRT in what follows.
@@ 3.1 Task definition
Extractive summarization aims at creating a summary $\mathcal{Y}$ from an input meeting $\mathcal{U}$, in which $\mathcal{U}$ consists of $|\mathcal{U}|$ elementary discourse units (EDUs) $[u_{1},u_{2},\dots,u_{|\mathcal{U}|}]$ and $\mathcal{Y}$ consists of $|\mathcal{Y}|$ EDUs, where $|\mathcal{Y}|\leq|\mathcal{U}|$ and $[y_{1},\dots,y_{|\mathcal{Y}|}]\in\mathcal{U}$. We borrow the notion of an EDU from SDRT, in which an EDU corresponds to the content of a single atomic clause so that the sentence “The remote should glow in the dark so that it’s easier to find” would be two EDUs [^1] We note that while the STAC and Molweni are segmented in this way, the transcripts of AMI and ICSI, which we use for our evaluations, are segmented slightly differently. The main difference is that when two consecutive clauses play the same discourse function in AMI and ICSI, they are not segmented. This means a sentence like “My name is Ed and I am the project manager” would not be segmented in these corpora because both independent clauses serve just to introduce the speaker. Because of the subtlety of this difference, we use the term EDU for simplicity..
It is worth to note that the input length of meeting $\mathcal{U}$ can be quite large and is often longer than the input size of most language models. E.g., the average length of ICSI meetings is 13317 tokens, while Llama’s (Touvron et al., 2023) context window is of 4096 tokens, which highlights the advantage of our graph-based method.
@@ 3.2 Discourse graphs
The graph for example 3.2333Example 3.2 is a slightly cleaned up version of an extract from meeting ES2011b of the AMI corpus. is given in Figure 2, where QAP stands for Question-Answer Pair and Ack is short for Acknowledgment.
@@ 4 Our model
To perform node-classification on a graph, we first need to provide an initial representation of the nodes. Since we are dealing with text, we simply use a text embedding model to get representations for EDUs. We chose MiniLM Reimers and Gurevych (2019), as it is a widely available sentence embedding model.
@@
To generate the graph structure, we need to add edges and edge labels for different discourse relations. Because discourse annotations require expert annotators and are very labor intensive, we opted to generate the edges and labels for our graphs with an automatic parser. In this paper, except for the ablation studies in which we consider the parsers of (Liu and Chen, 2021) and (Bennis et al., 2023), we use Deep Sequential Shi and Huang (2019).
@@
in which $\mathbf{N}_{i}^{r}$ denotes the set of neighbors of node $v_{i}$ under relation $r$ and $\mathbf{W}_{r}^{(l)}$ denotes learnable parameters at the $l$ layer for the relation $r$. Finally, we send the final representation $h_{i}^{l_{final}}$ to be classified by a sigmoid function on whether its content should be included or not in extractive summary.
@@ 5 Experiments
From the text summarization perspective (see Table 3), in addition to all the above baselines we evaluate: BERTSumExt (Liu and Lapata, 2019), which embeds individual input sentences with BERT and uses multiple inter-sentence transformer layers stacked on top of the BERT outputs to capture document-level features for extracting summaries; TextRank (Mihalcea and Tarau, 2004), a simple, graph-based and unsupervised extractive method; CoreRank (Tixier et al., 2017), whose approach relies on a Graph-of-Words representation of the meeting, graph degeneracy and submodularity for unsupervised extractive summarization; Finally, we use Lead-N as a heuristic metric, which takes the first N words of the meeting as its summary, N being the budgeted length of summaries.
@@ 6 Ablation studies
Figure 3: Frequency of the different discourse relation types in both AMI and ICSI.
@@ 6.1 Relation type ablation
As shown in Figure 3, the Deep Sequential parser predicts a similar distribution of relations for the AMI and ICSI datasets [^4] It is worth to note that certain relations such as Correction, Parallel, and Background are extremely rare in our graphs. This rarity is due to their infrequent occurrence in the STAC corpus, on which Deep Sequential was trained.. To evaluate the impact of each relation type on the RGCN classifier, we follow the same process as Feng et al. (2020) and train an RGCN that knows only one relation type at a time. That is, we feed a graph into an RGCN for which all edges are initialized with a value of 0, with the exception of edges that correspond to the target relation. Our observations indicate that within the AMI dataset, Correction, Acknowledgement, and Explanation play a significant role in classifying the EDUs, whereas for the ICSI dataset, the relations that stand out in terms of their influence on classification include Result, Contrast and Narration, and Explanation.
Figure 4: Impact of individual discourse relations on our RGCN classifier for both AMI and ICSI.
Figure 5: Impact of different rates of hidden relations on our RGCN classifier for both AMI and ICSI.
Figure 4 and 5 illustrate that for the AMI dataset, relation labels negatively impact classification. The highest F1 scores are achieved when no relations (None) are used in the RGCN, similar to the effect of using the rarely occurring Correction relation (see Figure 3), which is practically equivalent to not using any relationships. The RGCN also performs well when it only considers very rare relations, indicating that classification performance is driven by graph structure rather than relation types, except for Acknowledgments (the most frequent type in Figure 3) which, while being common usually point to small EDUs that seldom contribute to the summary. Conversely, for the ICSI dataset, the Result relation significantly influences classification compared to other relations, establishing the importance of certain relationships. Notably, in the ICSI dataset, omitting relations results in poorer classification outcomes, unlike with AMI. For both datasets, randomizing relationships among nodes worsens performance (see the rightmost column in Figure 4), indicating that nonsensical relations can hinder classification.
Figure 6: Impact of different rates of hidden edges on our GCN classifier for both AMI and ICSI.
Figure 7: Percentage of node in and out of the gold summary for the different values of node centrality.
@@ 6.2 Structural ablation
We conducted further experiments to investigate the impact of graph edges on the model’s overall classification performance. As shown in Figure 6, hiding edges decreases the performance of the classifiers for both AMI and ICSI. It appears that the graph structure generally has a more robust impact on classification than relation types do. Figure 7 shows that the more central a node is, the higher the chances of it being a part of the extractive summary is; in other words, the more common it is to refer to a certain EDU, the higher the odds of it being important.
@@ 6.3 Studying different parsers
As the previously described experiments were all carried out on graphs produced by Deep Sequential (Shi and Huang, 2019) only, we decided to also test the impact that different graph parsers have on our model. In this section, we compare the influence of graphs produced by Deep Sequential ($DS$) with that of graphs produced by Knowledge Enhanced ($KE$; Liu and Chen, 2021) and BERTLine ($BL$; Bennis et al., 2023). Results in Table 4 show that the three parsers lead to different classification results, $KE$ scoring higher in F1 score than the other two, and $DS$ having the highest precision score. This difference suggests that we should further investigate the nature of the differences between the predictions of these three models.
@@
In order to compare between the resulting graphs produced by different parsers, we employ the Weisfeiler-Lehman graph kernel (Shervashidze et al., 2011; Siglidis et al., 2020) to map these graphs into an embedding space in which structurally alike graphs are placed closer to each other, resulting in the following similarity scores: $sim(KE,DS)=0.3842$, $sim(KE,BL)=0.4810$, $sim(BL,DS)=0.7389$. These scores indicate varying levels of resemblance between the pairs, with the $BL$ and $DS$ showing the highest similarity.

### v3

@@ Abstract
We introduce an extractive summarization system for meetings that leverages discourse structure to better identify salient information from complex multi-party discussions. Using discourse graphs to represent semantic relations between the contents of utterances in a meeting, we train a GNN-based node classification model to select the most important utterances, which are then combined to create an extractive summary. Experimental results on AMI and ICSI demonstrate that our approach surpasses existing text-based and graph-based extractive summarization systems, as measured by both classification and summarization metrics. Additionally, we conduct ablation studies on discourse structure and relation type to provide insights for future NLP applications leveraging discourse analysis.
@@ 1 Introduction
In recent years, the task of meeting summarization has garnered significant attention Rennard et al. (2023b) due in part to the rising adoption of videoconferencing Kost (2020) and the accumulated amount of meeting recordings, highlighting the need for automatic, novel and efficient processing methods. Despite recent advances in natural language processing, the challenge of generating concise and coherent summaries from lengthy, multi-party meeting transcriptions remains. Large language models suffer from the middle curse (Liu et al., 2024) and struggle to use information in the middle of their context window, particularly for summarization (Ravaut et al., 2023).
@@
Our approach involves a novel combination of Graph Neural Networks (GNNs) and graph-based representations of discourse in which each node in a graph represents the content of an individual utterance and each edge represents a semantic relation between two utterances whose nature is specified by a label, e.g., Explanation, Correction, Question-Answer Pair or Acknowledgment. The task of extractive summarization is then couched as one of binary node-classification in which nodes are determined to be important or not, and the content of the nodes judged to be important is what determines the final extractive summary. This approach, in contrast to the generation at the whole graph-level (Feng et al., 2020), allows us to gain fine-grained insight into the interaction between the importance of an utterance in a conversation and its role in the overall discourse. A further advantage of a graph-based approach is that GNNs, whose attention is focused only on the neighbors of a given node, do not suffer from the context length restrictions imposed by the pairwise attentional computations of transformer-based models. This means that our extractive summarization approach can take an entire long meeting transcript as input without special treatment.
To validate our discourse-structure approach, we conduct extensive experiments on the AMI (Mccowan et al., 2005) and ICSI (Janin et al., 2003) corpora, comparing our method against diverse extractive summarization systems. The results demonstrate that our approach outperforms traditional techniques across evaluation metrics for both classification and summarization, including F1 score, ROUGE and BERTScore. Additionally, we present a detailed analysis of the impact of various graph construction strategies on summarization quality, offering insights into the mechanisms through which discourse structure influences content selection. Our study not only provides a novel methodological contribution to the field of automatic meeting summarization but also sheds light on the underlying discourse processes that shape effective summaries. Finally, SDRT based discourse structure has seldom been explored in the literature for its potential utility in downstream tasks. Our work is among the few studies that focus on this effort, especially for dialogues.
@@ 2 Related work
Most work focuses on abstractive summarization and tends to be organized in three interconnected streams (Rennard et al., 2023b). The first focuses on enhancing transcripts with additional information—such as annotations for dialogue acts (Goo and Chen, 2018), discourse structure (Feng et al., 2020) or visual cues (Liu et al., 2019)—that is assumed to be relevant for abstracting important content from a transcript. While adding information does show improvement, few papers actually quantify how much information is gained from adding linguistic features. Feng et al. (2020) used a global node to represent the discourse graph, aggregating all features into a single representative vector sent to an LSTM-based pointer decoder network for summarization; such an approach is not ideal for precisely evaluating the role of discourse structure in enhancing abstractive summaries.
The second stream aims to transform and compress meeting transcripts in order to produce cleaner and more condensed intermediate documents that can then be passed to downstream generation modules (Krishna et al., 2021; Oya et al., 2014). The third stream is dedicated to the development of multi-task systems, such as DialogLM (Zhong et al., 2022), a language model that is specifically designed to handle dialogue content. While we focus on extractive summarization, in exploiting discourse structure, our approach overlaps with work in the first stream above and lays groundwork for the second stream by providing a means of compressing meeting transcripts.
@@
Two primary theoretical frameworks have been developed to analyze complete discourse structures in the form of graphs, namely Rhetorical Structure Theory (RST; Mann and Thompson, 1987) and Segmented Discourse Representation Theory (SDRT; Asher, 1993; Lascarides and Asher, 2008). The framework of RST has been successfully applied for extractive summarization of documents (Marcu, 1997; Liu and Chen, 2019; Xu et al., 2020; Zhu et al., 2021; Bian et al., 2024) as well as abstractive summarization (Pu et al., 2023), however, these methods are not applicable to discussion by nature. However most work on SDRT discourse graphs has concentrated primarily on producing high-quality graphs (Shi and Huang, 2019; Liu and Chen, 2021; Bennis et al., 2023), with less emphasis on their applications to downstream tasks such as summarization, and only SDRT has been applied to build discourse structures for multi-party conversation (Asher et al., 2016), which is needed for meeting summarization.
A major limitation for research on discourse structure is of course the intense effort that goes into annotating datasets for discourse structure. There are only two readily available discourse-annotated corpora for multi-party conversation: STAC (Asher et al., 2016), a corpus of chats from an online version of the game Settlers of Catan, and Molweni (Li et al., 2020), an annotated version of the Ubuntu Chat Corpus. Both corpora are annotated in the style of SDRT.
@@ 3 Preliminaries
Due to its appropriateness for multi-party dialogue (Asher et al., 2016) and to the fact that it is the framework adopted for the STAC and Molweni corpora, we focus on SDRT in what follows.
@@ 3.1 Task definition
Extractive summarization aims at creating a summary $\mathcal{Y}$ from an input meeting $\mathcal{U}$, in which $\mathcal{U}$ consists of $|\mathcal{U}|$ elementary discourse units (EDUs) $[u_{1},u_{2},\dots,u_{|\mathcal{U}|}]$ and $\mathcal{Y}$ consists of $|\mathcal{Y}|$ EDUs, where $|\mathcal{Y}|\leq|\mathcal{U}|$ and $[y_{1},\dots,y_{|\mathcal{Y}|}]\in\mathcal{U}$. We borrow the notion of an EDU from SDRT, in which an EDU corresponds to the content of a single atomic clause so that the sentence “The remote should glow in the dark so that it’s easier to find” would be two EDUs (the second starting with ‘so that’) [^1] We note that while the STAC and Molweni are segmented in this way, the transcripts of AMI and ICSI, which we use for our evaluations, are segmented slightly differently. The main difference is that when two consecutive clauses play the same discourse function in AMI and ICSI, they are not segmented. This means a sentence like “My name is Ed and I am the project manager” would not be segmented in these corpora because both independent clauses serve just to introduce the speaker. Because of the subtlety of this difference, we use the term EDU for simplicity..
It is worth noting that the input length of meeting $\mathcal{U}$ will often be longer than the input size of most language models. The average length of ICSI meetings is 13317 tokens, for instance, while Llama’s (Touvron et al., 2023) context window is 4096 tokens, which highlights the advantage of our graph-based method.
@@ 3.2 Discourse graphs
The graph for example 3.2333Example 3.2 is a slightly cleaned up version of an extract from meeting ES2011b of the AMI corpus. is given in Figure 2.
@@ 4 Our model
To perform node-classification on a graph, we first need to provide an initial representation of the nodes. Since we are dealing with text, we simply use a text embedding model to get representations for EDUs. We chose MiniLM Reimers and Gurevych (2019), as it is a widely accessible, state-of-the-art model for sentence embeddings.
@@
To generate the graph structure, we need to add edges and edge labels for different discourse relations. Because discourse annotations require expert annotators and are very labor intensive, we opted to generate the edges and labels for our graphs with an automatic parser. In this paper, except for the ablation studies in which we consider the parsers of Liu and Chen (2021) and Bennis et al. (2023), we use Deep Sequential Shi and Huang (2019).
@@
in which $\mathbf{N}_{i}^{r}$ denotes the set of neighbors of node $v_{i}$ under relation $r$ and $\mathbf{W}_{r}^{(l)}$ denotes learnable parameters at the $l$ layer for the relation $r$. Finally, we send the final representation $h_{i}^{l_{final}}$ to be classified by a sigmoid function on whether its content should be included in the extractive summary.
@@ 5 Experiments
From the text summarization perspective (see Table 3), in addition to all the above baselines we evaluate: BERTSumExt (Liu and Lapata, 2019), which embeds individual input sentences with BERT and uses multiple inter-sentence transformer layers stacked on top of the BERT outputs to capture document-level features for extracting summaries; TextRank (Mihalcea and Tarau, 2004), a simple, graph-based and unsupervised extractive method; CoreRank (Tixier et al., 2017), whose approach relies on a Graph-of-Words representation of the meeting, graph degeneracy and submodularity for unsupervised extractive summarization. GPT-4, in which we ask to extract the most important utterances in a one-shot fashion to create the extractive summary (see Appendix A.1). Finally, we use Lead-N as a heuristic metric, which takes the first N words of the meeting as its summary, N being the budgeted length of summaries in our case.
@@ Results.
Method Gold RGCN Mixhop LM MLP CP
Gold 0 199 216 264 344 446
RGCN 201 0 196 267 355 438
Mixhop 184 204 0 257 351 392
GPT-4 136 133 143 0 305 -166
MLP 56 45 49 95 0 -1110
Table 4: Method ranking using LLM-as-a-Judge. Each row shows how many times a summary has been preferred over the model in the corresponding column. The Copeland Score is calculating by removing the number of losses to the number of wins of each models.
5.1 Extractive summaries ranking
Given the similarity in performance across methods, we conducted a ranking of the summaries using LLM-as-a-Judge (Zheng et al., 2024). We rank five generated summaries based on their informativeness and fluidity using GPT-4. In our case, we evaluate the summaries for each AMI meeting by ranking the outputs of the following methods: RGCN, MixHop GCN, GPT-4, MLP, and the human-created extractive summary. To carry out the ranking, we prompt GPT-4 with: “From this meeting: {Meeting}, rank these five summaries in order from best to worst based on how informative and readable they are {Summaries}. Only output the names of the summaries in order, and nothing else.” To ensure robustness and reduce variability, we repeat this question 20 times for each meeting. Results in Table 4 show that RGCN performs best, achieving the highest Copeland Score and consistently outscoring other methods in terms of informativeness and readability, showing the importance of taking relations into account. The fact that MixHop GCN ranks second suggests that graph-based models (RGCN, MixHop GCN) produce more effective extractive summaries than GPT-4 and MLP.
@@ 6 Ablation studies
Figure 3: Frequency of discourse relation types in AMI and ICSI predicted by Deep Sequential parser.
@@ 6.1 Relation type ablation
As shown in Figure 3, the Deep Sequential parser predicts a similar distribution of relations for the AMI and ICSI datasets [^4] It is worth noting that certain relations such as Correction, Parallel, and Background are extremely rare in our graphs. This rarity is due to their infrequent occurrence in the STAC corpus, on which Deep Sequential was trained.. To evaluate the impact of each relation type on the RGCN classifier, we follow the same process as Feng et al. (2020) and train an RGCN that knows only one relation type at a time. That is, we feed a graph into an RGCN for which all edges are initialized with a value of 0, with the exception of edges that correspond to the target relation. Our observations from Figure 4(a) indicate that within the AMI dataset, Correction, Acknowledgement, and Explanation play a significant role in classifying the EDUs, whereas for the ICSI dataset, the relations that stand out in terms of their influence on classification include Result, Contrast, Narration, and Explanation.
(a) Impact of individual discourse relations on our RGCN classifier for both AMI and ICSI.
(b) Impact of different rates of hidden relations on our RGCN classifier for both AMI and ICSI.
(c) Impact of different rates of hidden edges on our GCN classifier for both AMI and ICSI.
(d) Percentage of node in and out of the gold summary for the different values of node centrality.
Additionally, Figure 4(a) and 4(b) illustrate that for the AMI dataset, relation labels negatively impact classification. The highest F1 scores are achieved when no relations (None) are used in the RGCN, similar to the effect of using the rarely occurring Correction relation (see Figure 3), which is practically equivalent to not using any relations. The RGCN also performs well when it only considers very rare relations, indicating that classification performance is driven by graph structure rather than relation types, except for Acknowledgments (the most frequent type in Figure 3) which, while being common, usually point to small EDUs that seldom contribute to the summary. Conversely, for the ICSI dataset, the Result relation significantly influences classification compared to other relations, establishing the importance of certain relations. Notably, in the ICSI dataset, omitting relations results in poorer classification outcomes, unlike with AMI. For both datasets, randomizing relations among nodes worsens performance (see the rightmost column in Figure 4(a)), indicating that nonsensical relations can hinder classification. Synergistic role of different relations are discussed in Appendix B.
@@ 6.2 Structural ablation
We conducted further experiments to investigate the impact of graph edges on the model’s overall classification performance. As shown in Figure 4(c), hiding edges decreases the performance of the classifiers for both AMI and ICSI. It appears that the graph structure generally has a more robust impact on classification than relation types do. Figure 4(d) shows that the more central a node is, the higher the chances of it being a part of the extractive summary is; in other words, the more common it is to refer to a certain EDU, the higher the odds of it being important. Other ways to weigh the importance of EDUs such as length of utterances, speaker changes, length of links (how many utterances between two connected nodes) do not have impact on classification.
@@ 6.3 Studying different parsers
As the previously described experiments were all carried out on graphs produced by Deep Sequential (Shi and Huang, 2019) only, we decided to also test the impact that different graph parsers have on our model. In this section, we compare the influence of graphs produced by Deep Sequential ($DS$) with that of graphs produced by Knowledge Enhanced ($KE$; Liu and Chen, 2021) and BERTLine ($BL$; Bennis et al., 2023) as well as graphs generated by GPT-4 parser we develop in the Appendix A.2. Results in Table 5 show that the four parsers lead to heterogeneous results, $KE$ scoring higher in F1 score than the other three, and $DS$ having the highest precision score. This difference suggesting we should investigate the nature of the differences between the model predictions. Additionally, the GPT-4 parser did not deliver a significant improvement compared to what is typically observed in other NLP tasks, suggesting that discourse parsing remains a challenge for LLMs.